# FacLens: Transferable Probe for Foreseeing Non-Factuality in Large Language Models

### v3

@@ Abstract
Despite advancements in large language models (LLMs), non-factual responses remain prevalent. Unlike extensive studies on post-hoc detection of such responses, this work studies non-factuality prediction (NFP), aiming to predict whether an LLM will generate a non-factual response to a question before the generation process. Previous efforts on NFP have demonstrated LLMs’ awareness of their internal knowledge, but they still face challenges in efficiency and transferability. In this work, we propose a lightweight NFP model named Factuality Lens (FacLens), which effectively probes hidden representations of questions for the NFP task. Besides, we discover that hidden question representations sourced from different LLMs exhibit similar NFP patterns, which enables the transferability of FacLens across LLMs to reduce development costs. Extensive experiments highlight FacLens’s superiority in both effectiveness and efficiency.
@@ 1 Introduction
Large language models (LLMs) have shown impressive abilities in understanding and generating coherent text (OpenAI, 2023; Meta, 2024; Jiang et al., 2023), yet they usually struggle to provide accurate facts, leading to the generation of non-factual responses (Zhang et al., 2023; Cui et al., 2024). Extensive studies have devoted to detecting the non-factual responses, a task we name non-factuality detection (NFD) (Manakul et al., 2023; Azaria & Mitchell, 2023; Chen et al., 2024a; 2023; Min et al., 2023b). However, these post-hoc methods require response generation, which incurs computational overhead. Therefore, this paper studies non-factuality prediction (NFP), which predicts the likelihood of an LLM generating a non-factual response before the response generation. Figure 1 (a) illustrates the difference between NFD and NFP.
To address the NFP problem, researchers have proposed to extract and analyze specific tokens within a question (Mallen et al., 2023; Yüksekgönül et al., 2024), making their methods applicable to specific questions. For more general questions, some methods employ multi-round conversations with the LLM or fine-tune the entire LLM for NFP (Luo et al., 2023; Kadavath et al., 2022). While they have highlighted LLMs’ awareness of whether they possess certain knowledge, two limitations remain: (1) current NFP models lack a more efficient way to leverage such knowledge awareness, and (2) they are trained for individual LLMs, lacking transferability for rapid adaptation to new LLMs.
@@
Inspired by recent studies on monitoring and manipulating hidden representations to improve LLM performance (Zou et al., 2023; Zhang et al., 2024; Chen et al., 2024b), we assume that the knowledge awareness has been embedded in the hidden representation of a question. Based on this, we design a lightweight model, Factuality Lens (FacLens), which effectively probes hidden representations of input questions for NFP. Figure 1 (b) illustrates the workflow of FacLens. With its lightweight structure and rapid acquisition of hidden question representations, FacLens achieves high efficiency in both training and prediction (see Table 2). To obtain the training data of FacLens, we prompt the target LLM to produce responses to questions from high-quality question-answering (QA) datasets. We then compare the LLM-generated answers with the golden answers, assigning binary factual/non-factual labels to the responses. Nevertheless, extending FacLens to support multiple LLMs becomes resource-intensive and time-consuming, because each LLM must conduct response generation for the training data construction. Fortunately, we discover the transferability of FacLens, allowing us to assign the binary labels on just one LLM and apply unsupervised domain adaptation to rapidly apply FacLens to other LLMs without collecting new labels, thereby improving the development efficiency.
@@
The transferability of FacLens is inspired by research on human cognition, which shows individuals with similar cognitive styles and encoding strategies exhibit similar brain activity when performing the same task (Miller et al., 2012). We thus hypothesize that different LLMs have similar cognitive patterns in terms of knowledge awareness (i.e., similar NFP patterns), as current LLMs generally follow the Transformer (Vaswani et al., 2017) architecture and share overlapping training corpora. To validate this hypothesis, we define a collection of hidden question representations sourced from a certain LLM as a data domain. Our experimental results show that a FacLens trained on data from multiple domains performs similarly to the one trained on a single domain, suggesting that different individual domains do not have a significant concept shift (Moreno-Torres et al., 2012). [^1] Not having a significant concept shift implies highly consistent conditional distributions $P(\mathbf{y}|\mathbf{X})$ between data domains. Building on this insight, we can quickly apply FacLens to support a new LLM through unsupervised domain adaptation (Smola et al., 2007; Liu et al., 2022). To enhance the adaptation process, we introduce a question-aligned strategy to improve the efficacy of mini-batch training.
Figure 1: Illustrations of the objective and workflow of FacLens. We explore ante-hoc NFP by proposing a lightweight model named FacLens, which leverages hidden question representations to efficiently predict the likelihood of non-factual responses prior to their generation. Furthermore, we investigate the transferability of FacLens across multiple LLMs, enabling a cross-LLM FacLens to improve the overall development efficiency.
Overall, this paper makes the following contributions:
@@
  • Findings. We verify that hidden question representations in LLMs contain valuable information for NFP, i.e., LLMs’ activation during question-thinking mostly reveal whether they know the factual answers. Furthermore, we show that different LLMs exhibit similar NFP patterns to enable transfer learning of cross-LLM FacLens, which can align with human cognitive processes.
@@
  • Method. We propose a NFP model named FacLens, which features efficient development and application. Furthermore, we enable effective unsupervised domain adaptation of FacLens with a question-aligned mini-batch training strategy. To the best of our knowledge, this is a pioneer work to train a transferable NFP model for multiple LLMs.
@@
  • Experiments. All analyses are based on extensive experiments involving four popular open-source LLMs and three widely used factual QA datasets. The results show that FacLens outperforms the baselines in terms of both AUC metric and runtime. Human evaluation also demonstrates the effectiveness of FacLens.
@@ 2 Related Work
Hidden Representations in LLMs. The monitor and manipulation of hidden representations to improve LLM performance, known as representation engineering (RepE) (Zou et al., 2023), has been widely used to detect or control the factuality of LLM outputs. Studies like SAPLMA (Azaria & Mitchell, 2023) and MIND (Su et al., 2024) leverage hidden representations of LLM-generated responses for post-hoc NFD. KEEN Gottesman & Geva (2024) uses the hidden representations of input entities to regress the LLM’s accuracy on a QA task. TruthX (Zhang et al., 2024) edits the hidden representations of LLM-generated responses via an edit vector to activate the truthfulness of the LLM’s responses. Activation Decoding (Chen et al., 2024b) reveals that an LLM’s responses are closely tied to the representations of input entities. However, the effectiveness of hidden entity representations for NFP remains under-explored. In this paper, we show that question-level representations are more effective for NFP than entity-level representations.
Non-Factuality Prediction in LLMs. The entity popularity-based method (Mallen et al., 2023) assumes that LLMs are more familiar with questions about popular entities and uses Wikipedia page views to approximate the entity popularity. However, not every question contains entities that exactly match a Wikipedia entry. Recently, SAT Probe (Yüksekgönül et al., 2024) predicts based on the LLM’s attention to specific constraint tokens. The authors restrict the types and formats of questions to directly identify the constraint tokens. However, extracting constraint tokens from free-form questions can be challenging. Self-Familiarity (Luo et al., 2023) estimates an LLM’s familiarity with the requested entities through multi-round conversations with the LLM, requiring carefully-crafted prompts to engage the LLM multiple times. Besides, researchers fully fine-tune the LLM for NFP (Kadavath et al., 2022) (termed Self-Evaluation). Nevertheless, this incurs significant computational costs and can inhibit the LLM’s generalization ability (Yang et al., 2024b). In contrast, FacLens exhibits both good applicability and high efficiency. Moreover, this is a pioneer work that explores similar NFP patterns in hidden representations sourced from different LLMs.
Notably, a recent work (Liang et al., 2024) has demonstrated a correlation between hidden question representation and an LLM’s self-consistency (Manakul et al., 2023; Wang et al., 2023), where the self-consistency reflects the consistency of the LLM’s multiple responses to the same questions. However, self-consistency does not imply factuality. For example, LLMs may consistently provide an incorrect answer or refuse to respond with statements like, “I apologize, but I don’t have information on …” when unable to provide the factual answer.
@@ 3.1 Problem Definition
In this section, we formally define the concept of non-factual response and the problem of NFP for LLMs. Subsequently, we define the problem of transferable NFP across LLMs, which has rarely been discussed before.
@@
Non-Factual Response. Given an LLM $m\in\mathcal{M}$ and a question $q\in\mathcal{Q}$, $m$ generates an answer $s$ to the question. If the answer $s$ fails to convey the queried fact, it is a non-factual response.
@@
Non-Factuality Prediction in an LLM (NFP). Given an LLM $m\in\mathcal{M}$ and a question $q\in\mathcal{Q}$, the objective is to learn a function $f\left(m,q\right)\rightarrow y$, where $y=1$ if $m$ will generate a non-factual response to $q$ and $y=0$ otherwise.
@@
Transferable Non-Factuality Prediction Across LLMs. Given an LLM $m_{1}\in\mathcal{M}$, an LLM $m_{2}\in\mathcal{M}$, and a question set $\mathcal{Q}$, NFP labels have been collected based on $\mathcal{Q}_{train}\subset\mathcal{Q}$ for $m_{1}$, resulting in a training set $\{\left(\left(m_{1},q_{i}\right),y_{1,i}\right)\}_{q_{i}\in\mathcal{Q}_{train}}$. The goal is to utilize the training set and $m_{2}\in\mathcal{M}$ to learn a function $f\left(m,q\right)\rightarrow y$, where $m\in\{{m_{1},m_{2}}\}$ and $q\in\mathcal{Q}$.
3.2 Non-Factuality Prediction (NFP) Datasets
NFP Dataset Construction. Given an LLM $m$ and a QA dataset, for each question $q\in\mathcal{Q}$, we assign a binary label $y$ to the $\left(m,q\right)$ pair, where $y=1$ if $m$ fails to generate the golden answer for $q$, and $y=0$ otherwise. The goal of NFP is to predict the labels prior to answer generation. Specifically, we follow previous work (Mallen et al., 2023) to adopt QA datasets with short answers like entity mentions, and mark an LLM’s response as non-factual (i.e., $y=1$) if no sub-string of the response matches any of the gold answers. [^2] The annotation method ensures accurate labeling of all positive samples. We randomly sample 20 negative samples from each NFP dataset, deriving $20\times 12=240$ negative samples, and manually checked their labels’ quality. Given that all positive samples constitute 72.2% of the dataset, the correct label ratio is 97.0%. To ensure the experimental reproducibility, we set the LLM’s decoding strategy to greedy search rather than top-$p$ or top-$k$ sampling. We have also run the sampling-based decoding for response generation, and find that all the experimental conclusions in this paper still hold true. In this work, we consider four LLMs and three QA datasets, which results in $4\times 3=12$ NFP datasets. In each NFP dataset, consisting of samples in the form of $\left((m,q),y\right)$, we randomly sample 20% samples for training, 10% samples for validation, and use the remaining samples for testing.
LLMs & QA Datasets. We conduct experiments on four widely-used open-source LLMs: LLaMA2-7B-Chat (Touvron et al., 2023), LLaMA3-8B-Instruct (Meta, 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), and Qwen2-1.5B-Instruct (Yang et al., 2024a). These LLMs have been instruction-tuned for conversational engagement. We pose questions from three widely-used QA datasets: PopQA (PQ) (Mallen et al., 2023), Entity Questions (EQ) (Sciavolino et al., 2021), and Natural Questions (NQ) (Kwiatkowski et al., 2019). Detailed statistics of these QA datasets are provided in Appendix A.
@@ 4 Methodology
In this section, we propose a lightweight and transferable NFP model named FacLens. Our findings show that the hidden representations of users’ input questions in LLMs contain valuable patterns that are useful for NFP. Additionally, we show that hidden question representations derived from different LLMs share similar NFP patterns, facilitating the transfer learning of FacLens across multiple LLMs. Finally, we implement the cross-LLM FacLens using unsupervised domain adaptation.
@@ 4.1 FacLens
Inspired by representation engineering (Zou et al., 2023; Chuang et al., 2024; Li et al., 2023; Azaria & Mitchell, 2023; Su et al., 2024; Zhang et al., 2024; Yüksekgönül et al., 2024), we posit that before an LLM generates a response, the hidden question representation in the LLM contain potential patterns for predicting whether the LLM can generate the requested facts. In this section, we implement FacLens to verify this hypothesis.
@@
Factuality Lens (FacLens). We introduce FacLens, a learnable lightweight network to extract useful patterns from hidden question representations for NFP. Given an LLM $m$ and an input question $q$, we can quickly acquire the hidden states corresponding to input tokens. In a certain layer, we use the hidden states corresponding to the last input token as the question’s hidden representation $\mathbf{x}$. Then we use an encoder $g_{{enc}}$ to transform the question’s hidden representation into a latent feature space, where we presume that the NFP patterns are represented. Afterwards, a linear classifier $g_{{clf}}$ is set upon $g_{{enc}}$ for classification. Formally, based on the $\ell$-th hidden layer of $m$, FacLens predicts by,
$\mathbf{p}=f\left(m_{\leq\ell},q\right)=g_{{clf}}\left(g_{{enc}}\left(m_{\leq\ell}\left(q\right)\right)\right)=g_{{clf}}\left(g_{{enc}}\left(\mathbf{x}\right)\right)$ (1)
@@
where $m_{\leq\ell}\left(\cdot\right)$ denotes the function composed of the $\ell$-th transformer layer and its preceding layers, $g_{{enc}}$ is implemented by a lightweight multi-layer perceptron (MLP) [^3] We aim to verify that hidden question representations contain useful patterns for the NFP task. Exploring other model architectures for the NFP pattern extraction is beyond the scope of this paper., $g_{{clf}}$ is implemented by a linear layer with the Softmax function, and $\mathbf{p}$ is a two-dimensional vector revealing the probability of (not) producing non-factual responses. Based on a set of labeled NFP instances $\{\left(m_{\leq\ell}\left(q_{i}\right),y_{i}\right)\}_{q_{i}\in\mathcal{Q}_{train}\cup\mathcal{Q}_{val}}$, where $Q_{train}$ and $Q_{val}$ denote question sets used for training and validation, respectively, we can train a FacLens for $m$ with the classic cross-entropy (CE) loss.
Hidden Question Representation vs. Hidden Response Representation. In order to explore the feasibility of using only hidden representations of users’ input questions for non-factuality identification, we compare FacLens with SAPLMA (Azaria & Mitchell, 2023) and INSIDE (Chen et al., 2024a). They are two typical post-hoc methods that employ hidden representations of LLM-generated responses. SAPLMA (Azaria & Mitchell, 2023) is a classifier trained for NFD based on the hidden response representations. INSIDE (Chen et al., 2024a), on the other hand, leverages the eigenvalues of the covariance matrix of responses’ representations to measure self-consistency, assuming that inconsistent responses to the same question tend to be unreliable. As shown in Figure 2, FacLens stands out as a good choice. Specifically, INSIDE adopt the self-consistency assumption. However, self-consistency does not imply factuality. For instance, LLMs may consistently produce incorrect answers or decline to respond with statements like, “I apologize, but I don’t have information on …,” when they cannot provide a factual answer. In contrast, SAPLMA and FacLens exhibit competitive performance. Similar to FacLens, SAPLMA prefers hidden representations at the middle layer. Therefore, we set both FacLens and SAPLMA to use the hidden representations from the middle layer. We observe that FacLens outperforms SAPLMA on most NFP datasets. That is because responses generated by an LLM could contain some interference information, so the corresponding hidden representations may be sub-optimal for NFD. For example, a user inputs a question “What is the occupation of Taylor Swift?”, an LLM responds with “Taylor Swift is an American singer-songwriter and pop culture icon, born on December 13th, 1989. In addition to her music career, Swift is also involved in philanthropy and activism, advocating for various causes.” which correctly answers the question but includes information that is irrelevant to the question.
Figure 2: Comparison between hidden question representations and hidden response representations for identifying non-factual responses. FacLens is our ante-hoc method that employs hidden question representations, while INSIDE and SAPLMA are two typical post-hoc methods that employ hidden response representations.

### v4

# FacLens: Transferable Probe for Foreseeing Non-Factuality in Fact-Seeking Question Answering of Large Language Models

@@ Abstract
Despite advancements in large language models (LLMs), non-factual responses still persist in fact-seeking question answering. Unlike extensive studies on post-hoc detection of these responses, this work studies non-factuality prediction (NFP), predicting whether an LLM will generate a non-factual response prior to the response generation. Previous NFP methods have shown LLMs’ awareness of their knowledge, but they face challenges in terms of efficiency and transferability. In this work, we propose a lightweight model named Factuality Lens (FacLens), which effectively probes hidden representations of fact-seeking questions for the NFP task. Moreover, we discover that hidden question representations sourced from different LLMs exhibit similar NFP patterns, enabling the transferability of FacLens across different LLMs to reduce development costs. Extensive experiments highlight FacLens’s superiority in both effectiveness and efficiency.  [^1] Code: https://github.com/wyl7/FacLens.
FacLens: Transferable Probe for Foreseeing Non-Factuality in Fact-Seeking Question Answering of Large Language Models
Yanling Wang1thanks: Work done while at Zhongguancun Laboratory., Haoyang Li3, Hao Zou2, Jing Zhang3thanks: Corresponding authors., Xinlei He4, Qi Li2,522footnotemark: 2, Ke Xu2,5 1Zhipu AI  2Zhongguancun Laboratory  3Renmin University of China 4Hong Kong University of Science and Technology (Guangzhou)  5Tsinghua University yanlingwang777@gmail.comzhang-jing@ruc.edu.cnqli01@tsinghua.edu.cn
@@ 1 Introduction
Large language models (LLMs) have shown impressive abilities in understanding and generating coherent text (OpenAI, 2023; Meta, 2024; Jiang et al., 2023), yet they may provide non-factual responses in fact-seeking question answering (fact-seeking QA) (Zhang et al., 2023; Cui et al., 2024). Extensive studies have been devoted to detecting the non-factual responses, a task we name non-factuality detection (NFD) (Manakul et al., 2023; Azaria and Mitchell, 2023; Chen et al., 2024a, 2023; Min et al., 2023a). However, these post-hoc methods require response generation, which incurs significant computational overhead. Therefore, this paper studies non-factuality prediction (NFP), which predicts the likelihood of an LLM generating a non-factual response to a fact-seeking question before the response generation. Figure 1 (a) illustrates the difference between NFD and NFP.
Figure 1: Illustrations of the objective and workflow of FacLens for efficient and transferable ante-hoc NFP.
To solve the NFP problem, researchers have proposed predicting by analyzing specific tokens in a question (Mallen et al., 2023; Yüksekgönül et al., 2024), making these methods applicable to specific types of questions. For more general questions, approaches based on prompting or fine-tuning the LLM for NFP have been proposed Luo et al. (2023); Kadavath et al. (2022). However, two limitations persist: (1) current NFP models can be improved in terms of effectiveness and efficiency, and (2) they are designed for an individual LLM, lacking the transferability for rapid adaptation to new LLMs.
@@
Inspired by studies that monitor and manipulate hidden representations to improve LLMs’ performance (Zou et al., 2023; Zhang et al., 2024; Chen et al., 2024b), we hypothesize that knowledge awareness has been embedded in the hidden representations of fact-seeking questions. To investigate this, we propose a lightweight model, Factuality Lens (FacLens), and demonstrate its ability to probe the hidden question representations for NFP. Figure 1 (b) shows the workflow of FacLens. As hidden question representations can be efficiently obtained, and the model structure is lightweight, FacLens achieves high efficiency in both training and prediction (see Table 2). This makes it highly suitable for real-world LLM applications, helping to reduce non-factual responses to end-users while maintaining a minimal budget and latency. To construct the training data of FacLens, we prompt the target LLM to produce responses to questions from high-quality fact-seeking datasets. We then compare the LLM-generated answers with the golden answers, assigning binary factual/non-factual labels to the responses. Nevertheless, extending FacLens to support multiple LLMs becomes resource-intensive and time-consuming, because each LLM must conduct response generation for the training data construction. Fortunately, we discover the transferability of FacLens across different LLMs, allowing us to assign the binary labels for just one LLM and apply unsupervised domain adaptation (DA) to rapidly apply FacLens to other LLMs without collecting new labels, thereby improving the efficiency in development.
@@
Investigating the transferability of FacLens is inspired by the research on human cognition (Miller et al., 2012), which shows that individuals with similar cognitive styles and encoding strategies exhibit similar brain activity when performing the same task. Given current LLMs generally follow the Transformer (Vaswani et al., 2017) architecture and share overlapping training corpora Gao et al. (2021); com (2024); Kocetkov et al. (2023), we hypothesize that different LLMs have similar cognitive patterns in terms of knowledge awareness (i.e., similar NFP patterns). To validate this hypothesis, we define a collection of hidden question representations sourced from a certain LLM as a data domain. Our experiments demonstrate that a FacLens trained on data from multiple domains performs similarly to the one trained on a single domain, suggesting that different individual domains do not have a significant concept shift (Moreno-Torres et al., 2012). [^2] Not having a significant concept shift implies highly consistent conditional distributions $P(\mathbf{y}|\mathbf{X})$ between domains. Building on this insight, we can quickly apply a trained FacLens to a new LLM through unsupervised domain adaptation (DA) (Smola et al., 2007; Liu et al., 2022). Additionally, we propose a question-aligned strategy to enhance the mini-batch-based DA.
This paper makes the following contributions:
@@
  • Findings. We verify that hidden question representations in an LLM contain valuable information for NFP in fact-seeking QA, i.e., LLMs’ activation during question-embedding mostly reveals whether they know the facts. Moreover, we show that different LLMs exhibit similar NFP patterns to enable a cross-LLM NFP model.
@@
  • Method. We propose a lightweight and transferable NFP model named FacLens, enabling efficient development and application. To our knowledge, this is a pioneer work to train a transferable NFP model for multiple LLMs.
@@
  • Experiments. We conduct extensive experiments, involving four open-source LLMs and three widely used fact-seeking QA datasets. The results show that FacLens outperforms baselines in terms of AUC metric and runtime.
@@ 2 Related Work
Hidden Representations in LLMs. Hidden representations in LLMs have been shown to encode valuable information that can be leveraged for various tasks Zou et al. (2023). In terms of LLMs’ factuality, studies like SAPLMA Azaria and Mitchell (2023) and MIND Su et al. (2024) leverage hidden representations of LLM-generated responses for post-hoc NFD. TruthX Zhang et al. (2024) edits hidden representations of LLM-generated responses via an edit vector to enhance the LLM’s truthfulness. Activation Decoding Chen et al. (2024b) reveals that an LLM’s responses are closely tied to the representations of input entities but do not validate their effectiveness in NFP.
Non-Factuality Prediction in LLMs. We categorize the studies into token-based and non-token-based methods. The entity popularity-based method Mallen et al. (2023) focuses on input entity tokens, assuming that LLMs are more familiar with questions about popular entities and estimating entity popularity based on Wikipedia page views. However, not every question contains entities that exactly match a Wikipedia entry. SAT Probe Yüksekgönül et al. (2024) predicts based on the LLM’s attention to specific constraint tokens. The authors restrict the types and formats of questions to facilitate the identification of the constraint tokens. However, extracting constraint tokens from free-form questions is non-trivial. Without focusing on specific tokens, Self-Familiarity Luo et al. (2023) estimates an LLM’s familiarity with the requested facts through multi-round conversations with the LLM, requiring carefully crafted prompts to engage the LLM multiple times, resulting in low prediction efficiency. Alternatively, researchers fully fine-tune the LLM for NFP Kadavath et al. (2022) (termed Self-Evaluation). However, this approach incurs significant computational costs and may hinder the LLM’s generalization ability Yang et al. (2024b). In contrast, FacLens exhibits good applicability, high efficiency, and enables cross-LLM domain adaptation.  [^3] Hidden question representation has been utilized for predicting an LLM’s self-consistency Liang et al. (2024). Notably, self-consistency does not equate to factuality; for instance, an LLM can consistently produce incorrect answers.
@@ 3.1 Problem Definition
Non-Factual Response. Given an LLM $m\in\mathcal{M}$ and a fact-seeking question $q\in\mathcal{Q}$, $m$ generates an answer $s$. If the answer $s$ fails to convey the queried fact, it is a non-factual response.
@@
Non-Factuality Prediction in an LLM (NFP). Given an LLM $m\in\mathcal{M}$ and a fact-seeking question $q\in\mathcal{Q}$, the objective is to learn a function $f\left(m,q\right)\rightarrow y$, where $y=1$ if $m$ will generate a non-factual response to $q$ and $y=0$ otherwise.
@@
Transferable Cross-LLM NFP. Given LLMs $m_{1},m_{2}\in\mathcal{M}$ and a fact-seeking question set $\mathcal{Q}$, NFP labels have been constructed based on $\mathcal{Q}_{train}\subset\mathcal{Q}$ for $m_{1}$, deriving a training set $\{\left(\left(m_{1},q_{i}\right),y_{1,i}\right)\}_{q_{i}\in\mathcal{Q}_{train}}$. The goal is to utilize the training set and $m_{2}$ to learn a function $f\left(m,q\right)\rightarrow y$, where $m\in\{{m_{1},m_{2}}\}$ and $q\in\mathcal{Q}$.
3.2 NFP Datasets
Dataset Construction. Given an LLM $m$ and a fact-seeking QA dataset, for each question $q\in\mathcal{Q}$, we assign a binary label $y$ to the $\left(m,q\right)$ pair, where $y=1$ if $m$ fails to generate the golden answer for $q$, and $y=0$ otherwise. Our goal is to predict the labels prior to answer generation. Notably, a fact-seeking question asks for objective and verifiable information, such as dates, locations, and entities. Examples include “In which year was the Eiffel Tower built?” and “Which city is the capital of France?”. Their answers are naturally short and precise. Therefore, we follow previous work (Mallen et al., 2023) to mark an LLM’s response as non-factual (i.e., $y=1$) if no sub-string of the response matches any of the gold answers. [^4] The labeling method ensures accurate labels of all positive samples. We randomly sample 20 negative samples from each NFP dataset, deriving $20\times 12=240$ negative samples, and manually checked their labels’ quality. Given that all positive samples constitute 72.2% of the dataset, the ratio of correct labels is 97.0%. We consider four LLMs and three QA datasets in the main body of the paper, deriving $4\times 3=12$ NFP datasets. In each NFP dataset, consisting of samples in the form of $\left((m,q),y\right)$, we randomly sample 20% data for training, 10% data for validation, and use the remaining data for testing.
LLMs & QA Datasets. LLaMA2-7B-Chat (Touvron et al., 2023), LLaMA3-8B-Instruct (Meta, 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), and Qwen2-1.5B-Instruct (Yang et al., 2024a) are used for experiments. These LLMs have been instruction-tuned for conversational engagement. We pose questions from three widely-used QA datasets: PopQA (PQ) (Mallen et al., 2023), Entity Questions (EQ) (Sciavolino et al., 2021), and Natural Questions (NQ) (Kwiatkowski et al., 2019). Detailed statistics of these datasets are provided in Appendix A. To ensure reproducibility, we set each LLM’s decoding strategy to greedy search rather than top-$p$ or top-$k$ sampling. We have also run the sampling-based decoding, and find that the paper’s experimental conclusions still hold true.
@@ 4.1 FacLens
Given an LLM $m$ and a fact-seeking question $q$, we can quickly acquire the hidden states corresponding to input tokens. In a certain layer, we use the hidden states corresponding to the last input token as the question’s hidden representation $\mathbf{x}$. Then we use an encoder $g_{{enc}}$ to transform the question’s hidden representation into a latent feature space, where we presume that the NFP patterns are represented. Afterwards, a linear classifier $g_{{clf}}$ is set upon $g_{{enc}}$ for classification. Formally, based on the $\ell$-th hidden layer of $m$, FacLens predicts by,
$\mathbf{p}=g_{{clf}}\left(g_{{enc}}\left(m_{\leq\ell}\left(q\right)\right)\right)=g_{{clf}}\left(g_{{enc}}\left(\mathbf{x}\right)\right)$ (1)
@@
where $m_{\leq\ell}\left(\cdot\right)$ denotes the function composed of the $\ell$-th transformer layer and its preceding layers, $g_{{enc}}$ is implemented by a lightweight multi-layer perceptron (MLP) [^5] Our goal is to verify that hidden question representations contain useful patterns for the NFP task. Exploring other model architectures for the NFP pattern extraction is beyond the scope of this paper., $g_{{clf}}$ is implemented by a linear layer with the Softmax function, and $\mathbf{p}$ is a two-dimensional vector revealing the probability of (not) producing non-factual responses. Based on a set of labeled NFP instances $\{\left(m_{\leq\ell}\left(q_{i}\right),y_{i}\right)\}_{q_{i}\in\mathcal{Q}_{train}\cup\mathcal{Q}_{val}}$, where $Q_{train}$ and $Q_{val}$ denote question sets used for training and validation, respectively, we can train a FacLens for $m$ with the classic cross-entropy (CE) loss.
Figure 2: Performance comparison of FacLens (ante-hoc NFP) with INSIDE and SAPLMA (post-hoc NFD). Trends are consistent on EQ and NQ (see Appendix D.1).
Observation: While post-hoc NFD methods leverage more information, the ante-hoc FacLens has the potential to perform better. In Figure 2, we compare FacLens with SAPLMA (Azaria and Mitchell, 2023) and INSIDE (Chen et al., 2024a), two representative post-hoc methods that leverage hidden representations of LLM-generated responses to identify non-factual responses. While post-hoc methods leverage more information (i.e., the LLM-generated responses), FacLens shows comparable and even better performance.